OpenStudy (anonymous):

If = 27° and x = 9 cm, what is the value of z to the nearest tenth of a centimeter? A) 19.8 B) 18.8 C) 10.1 D) 17.7

OpenStudy (bloomlocke367):

where's z?

OpenStudy (anonymous):

On the opposite side of Y

OpenStudy (bloomlocke367):

and y=27?

OpenStudy (anonymous):

No

OpenStudy (bloomlocke367):

what =27?

OpenStudy (anonymous):

I think its sin or something like that im not sure the symbol won't show up its like a little circle with a line through it

OpenStudy (bloomlocke367):



OpenStudy (anonymous):

Yes and in that angle is looks like this 

OpenStudy (bloomlocke367):

yea cx have you heard of SOHCAHTOA?

OpenStudy (anonymous):

No

OpenStudy (bloomlocke367):

well it's how most people remember whether to use sine, cosine or tangent.

OpenStudy (anonymous):

so whats the answer?

OpenStudy (bloomlocke367):

I can't just give you the answer, it's against the rules

OpenStudy (anonymous):

I only have 3 more min though

OpenStudy (bloomlocke367):

anyways, sin=opp/hyp cos=adj/hyp and tan=opp/adj

OpenStudy (bloomlocke367):

you have the opposite and are looking for the hypotenuse so you need to use sin..

OpenStudy (anonymous):

I don't have a caculator

OpenStudy (bloomlocke367):

do you have a graphing calculator? I don't have one, but you need one...

OpenStudy (bloomlocke367):

ohhhh

OpenStudy (bloomlocke367):

well sin(27)=0.4539904997

OpenStudy (bloomlocke367):

so you have 0.4539904997=9/?

OpenStudy (bloomlocke367):

then you have ?*0.4539904997=9

OpenStudy (bloomlocke367):

9/0.4539904997=19.8242...

OpenStudy (bloomlocke367):

there you go...

OpenStudy (bloomlocke367):

do you understand what I did?